# Agentic Large Language Models, a survey

### v2

@@ 4.3 Simulating Open-ended Societies
Agentic LLMs have enhanced abilities for perception, memory, reasoning, decision-making, and adaptive learning. They can display heterogeneous personality profiles (Gao et al., 2023a, 2024). Such features make them also suitable for interacting in open-ended multi-agent simulations without prior role assignment. This allows the study of emergent phenomena such as self-organizing behaviors, collective intelligence and the development of social conventions and norms. Being able to simulate such phenomena more realistically, using heterogeneous agents that communicate in natural language, meets long-standing interests from the social sciences. The structure of LLM-based agents suitable for such simulations is illustrated in Figure 24 and Figure 25.
4.3.1 Simulacra and Societies
@@
Park et al. (2023) introduced Generative Agents, an environment where users can interact with a simulated town populated by 25 LLM-based agents. Based on social simulacra techniques proposed earlier (Park et al., 2022), each agent was initiated with a unique persona and memory. For each agent, a record is kept of all the experiences and conversations in the simulation, used to synthesize higher-level reflections and plan behavior. The agents behave somewhat like characters in The Sims: they initiate conversations, form relationships, spread information, and coordinate impromptu group activities. Figure 26 depicts the agent architecture and Figure 27 shows an illustration of a simulation. The interactions are influenced by user input and are therefore semi-autonomous. This is illustrated by the example of a Valentine’s Day party: while multiple agents spread invitations to one another and show up at the right time with coordinated plans, the plan for the party was initiated with a user prompt. The agents developed believable social routines (such as daily schedules, and gossip) and even exhibit human-like character traits (some agents demonstrated deception or stubbornness, while others showed cooperation). These results show that social patterns can emerge from dynamic LLM interactions.
@@
AgentSociety is a simulation at a larger scale, involving over 10,000 agents (Piao et al., 2025b, a). It aims not only to study everyday social dynamics, but it also offers a testbed for computational social experiments. The authors discuss case studies of polarization, the spread of inflammatory messages, the effects of universal basic income policies, and the impact of external shocks such as hurricanes. Li et al. (2024c) also study the spread of misinformation using LLM agents. Their agents exhibit diverse profiles in terms of gender, age, and the Big Five personality traits. One of the findings is that encouraging comments does not significantly reduce the spread of misinformation, whereas publicly labeling information with accuracy scores and blocking specific influencers proved to be effective strategies, particularly in scale-free networks.
@@
AgentVerse is a multi agent system to study group dynamics (Chen et al., 2023c). Inspired by human group dynamics, it studies whether a group of expert agents can be more than the sum of its parts. Experiments on text understanding, reasoning, coding, tool utilization, and embodied AI confirm the effectiveness. Problem solving is split into four stages: (1) expert recruitment, (2) collaborative decision making, (3) action execution, and (4) evaluation, where, if the current state is unsatisfactory, a new iteration of the process is started for refinement (see Figure 28). Interestingly, agents manifest emergent behaviors such as volunteering, characterized by agents offering assistance to peers, or conformity, where agents adjust deviated behaviors to align with the common goal under the critics from others. Destructive behaviors were also observed, occasionally leading to undesired and detrimental outcomes.
@@
OASIS is a scalable social media simulator for Twitter/X and Reddit (Yang et al., 2024e). It supports modeling of up to one million LLM-agents. It is built on CAMEL and has role-based agents as its starting point. However, at its large scale, OASIS shows various social group phenomena, including spreading of (mis)information, group polarization, and herd effects. OASIS is built upon an Environment Server, Recommender System, Agent Module, Time Engine, and Scalable Inferencer (see Figure 29).
@@
Research predating LLMs already shows that agent societies may create an automated curriculum of ever increasing difficulty (Elman, 1991; Bengio et al., 2009; Silver et al., 2017; Soviany et al., 2022), requiring increasing levels of intelligent behavior from the agents (Racaniere et al., 2019). Similar results have been observed for LLMs (Feng et al., 2023). WebArena is an environment developed to study self-evolving curricula (Qi et al., 2024), which can also help robot training (Ryu et al., 2024) or to mitigate hallucination (Zhao et al., 2024b).
4.3.2 Emergent Social Norms
@@
Social norms play an important role in the predictability of individuals in groups (Axelrod, 1981, 1986). Cultural evolution studies how norms evolve at a society level when individuals transmit behavior through imitation, communication, and education (Boyd and Richerson, 1988). LLMs endow agents with the ability to communicate in natural language and have created more opportunities for multi-agent research into societies and the emergence of conventions and norms. Extensive overviews of such new possibilities are provided in (Mou et al., 2024; Savarimuthu et al., 2024; Xi et al., 2023). We discuss some of the new approaches in more detail.
@@
EvolutionaryAgent (Li et al., 2024b) studies agent alignment in a multi-agent system, with evolutionary methods that go beyond Reinforcement Learning from Human Feedback (see Figure 30). In the context of agent alignment to norms, the approach is controlled: it does not permit the evolution of social norms to be disorderly or random, but it also does not intervene in each step of their evolution. The authors define the initial social norms and a desired direction of evolution. Agents with higher fitness (more norm-conforming) are more likely to reproduce, leading to the diffusion of their strategies, gradually stabilizing and forming new social norms.
@@
Defining a complete, realistic, and complex virtual society is challenging. The purpose of the work is to study how, if such a virtual society existed, a system could further enable evolving intricate evolutionary behaviors of agents, and how this would lead to the emergence of new dynamics. The system provides provides a sandbox for investigating the safety of AI systems before they impact the real world.
@@
A different approach is based on Steels (1995)’s naming game, implemented with agents powered by LLaMa 3 and Claude 3.5 (Kouwenhoven et al., 2024; Ashery et al., 2024; Baronchelli, 2023). They find that globally accepted conventions or norms can spontaneously arise from local interactions between communicating LLMs. The authors also demonstrate how strong collective biases can emerge during this process, even when individual agents appear to be unbiased, and how minority groups of committed LLMs can drive social change by establishing new social conventions that can overturn established behaviors.
@@
The emergence of norms is studied at another level by Horiguchi et al. (2024). They explore the potential for LLM agents to spontaneously generate and adhere to normative strategies, building upon the foundational work of Axelrod’s metanorm games. Metanorms are norms enforcing the punishment of those who do not punish agents that are breaking norms (Axelrod, 1986). Controlling for personality traits vengefulness and boldness, they find that through dialogue, LLM agents can form complex social norms, metanorms, purely through natural language interaction. A related study evaluates the capability of LLMs to detect norm violations (He et al., 2024). Based on simulated data from 80 stories in a household context, they investigated whether 10 norms are violated, and found ChatGPT-4 being able for detect norm violations, with Mistral some distance behind.
@@
Qiu et al. (2024a) go a step beyond norms, and study the cultural and social awareness of LLM agents. They introduce CASA, a benchmark designed to assess LLM agents’ sensitivity to cultural and social norms across two web-based tasks: online shopping and social discussion forums. (CASA is based on WebArena (Qi et al., 2024).) Current LLMs perform significantly better in non-agent than in web-based agent environments, with agents achieving less than 10% awareness coverage and over 40% violation rates. However, using prompting and finetuning on specific datasets, cultural and social awareness can be improved.
@@
Inspired by Society of Mind (Minsky, 1988), cooperation mechanisms are explored in Zhang et al. (2023a)’s agentic LLM simulation. This simulation consists of four unique societies of LLM agents, where each agent is characterized by a specific trait (easy-going or overconfident) and engages in cooperation with a distinct thinking pattern (debate or reflection). They find that LLM agents show human-like social behaviors, such as conformity and consensus reaching, mirroring foundational social psychology theories. Figure 31 shows societies with different types of agents.
@@
The question whether groups of LLM agents can successfully engage in cross-national collaboration and debate is studied by (Baltaji et al., 2024). Multi agent discussions can support collective AI decisions that reflect diverse perspectives, although agents are susceptible to conformity due to perceived peer pressure. They can also lose track of their personas and opinions. Instructions that encourage debate increase the risk of errors.
4.3.3 Open-World Agents
@@
An important driver of agentic LLM research is the problem of plateauing LLM performance due to limited training data. Open World multi-agent interaction aims to address this problem, generating new interaction data with multi agent simulations. Machine learning can learn no more complexity than what is present in the dataset (or environment). The idea of an open world-model is that it can create infinite datasets or environments, in which agents can continue to learn, to keep improving their intelligence. How should such unlimited challenges be created? The advent of LLMs has given a new impulse to this research question: LLMs are used to solve an LLM-generated problem. This idea is followed, for example, in the multi-agent finetuning approach (Subramaniam et al., 2025).
@@
Current agents are primarily created and tested in simplified synthetic environments, leading to a disconnect with real-world scenarios. Zhou et al. (2023a) build an environment, Webarena, that is more realistic and reproducible. WebArena is an environment with fully functional websites from four common domains: e-commerce, social forum discussions, collaborative software development, and content management.
@@
Games are eminently suited as open-ended benchmarks for interactive behavior. Real-world tasks require handling intricate interactions, advanced spatial reasoning, long-term planning, and continuous exploration of new strategies. Balrog (Paglieri et al., 2024) incorporates reinforcement learning environments of varying levels of difficulty, ranging from tasks that are solvable by non-expert humans in seconds to challenging ones that may take years to master (such as the NetHack Learning Environment). They find that while current models achieve partial success in the easier games, they struggle significantly with more challenging tasks such as vision-based decision-making.
@@
Progress in machine learning depends on benchmark availability. As models evolve, there is a need to create benchmarks that can measure progress on new generative capabilities (Butt et al., 2024). BenchAgents decomposes the benchmark creation process into planning, generation, data verification, and evaluation, each of which is executed by an LLM agent. These agents interact with each other and utilize human-in-the-loop feedback to explicitly improve and flexibly control data diversity and quality. BenchAgents creates benchmarks to evaluate capabilities related to planning and constraint satisfaction.
@@
AgentBoard provides an evaluation of the breadth of existing benchmarks (Ma et al., 2024a). Benchmarks should have task diversity. It is necessary to cover various agent tasks such as embodied, web, and tool tasks. Additionally, multi round interaction is important, to mimic realistic scenarios. Existing benchmarks typically adopt single-round tasks. Furthermore, agents should be evaluated in partially-observable environments, to test if they can actively explore their surroundings. Existing agent benchmarks fail to satisfy all of these criteria (Ma et al., 2024a).
4.4 Discussion
@@ 4.4.1 Interaction Studies
When interacting in multi-agent environments, agentic LLMs show varying levels of performance on games with strategic and coordinated behavior. Enhancing models with reasoning capacities boosts performance, as is evidenced by GPT-o1’s better overall performance and the positive effect of adding probabilistic graphs (Xu et al., 2024a). Pre-defined roles and interaction protocols (cooperative or adversarial) help structure the communication between LLM agents while improving task performance. Role-playing frameworks, AI feedback loops, and debate moderation suggests that carefully coordinating multiple LLMs can harness their collective intelligence and yield outcomes that surpass single model performance.
@@
4.4.2 Use Cases: Emergent Behavior and Training Data
Emergent Behavior, Cooperation

### v3

@@ 4.3. Simulating Open-ended Societies
Agentic LLMs have enhanced abilities for perception, memory, reasoning, decision-making, and adaptive learning. They can display heterogeneous personality profiles (gao2023s; gao2024large). Such features make them also suitable for interacting in open-ended multi-agent simulations without prior role assignment. This allows the study of emergent phenomena such as self-organizing behaviors, collective intelligence and the development of social conventions and norms. Being able to simulate such phenomena more realistically, using heterogeneous agents that communicate in natural language, meets long-standing interests from the social sciences. The structure of LLM-based agents suitable for such simulations is illustrated in Figure 24 and Figure 25.
4.3.1. Simulacra and Societies
@@
park2023generative introduced Generative Agents, an environment where users can interact with a simulated town populated by 25 LLM-based agents. Based on social simulacra techniques proposed earlier (park2022social), each agent was initiated with a unique persona and memory. For each agent, a record is kept of all the experiences and conversations in the simulation, used to synthesize higher-level reflections and plan behavior. The agents behave somewhat like characters in The Sims: they initiate conversations, form relationships, spread information, and coordinate impromptu group activities. Figure 26 depicts the agent architecture and Figure 27 shows an illustration of a simulation. The interactions are influenced by user input and are therefore semi-autonomous. This is illustrated by the example of a Valentine’s Day party: while multiple agents spread invitations to one another and show up at the right time with coordinated plans, the plan for the party was initiated with a user prompt. The agents developed believable social routines (such as daily schedules, and gossip) and even exhibit human-like character traits (some agents demonstrated deception or stubbornness, while others showed cooperation). These results show that social patterns can emerge from dynamic LLM interactions.
@@
AgentSociety is a simulation at a larger scale, involving over 10,000 agents (piao2025agentsociety; piao2025polarization). It aims not only to study everyday social dynamics, but it also offers a testbed for computational social experiments. The authors discuss case studies of polarization, the spread of inflammatory messages, the effects of universal basic income policies, and the impact of external shocks such as hurricanes. li2024large also study the spread of misinformation using LLM agents. Their agents exhibit diverse profiles in terms of gender, age, and the Big Five personality traits. One of the findings is that encouraging comments does not significantly reduce the spread of misinformation, whereas publicly labeling information with accuracy scores and blocking specific influencers proved to be effective strategies, particularly in scale-free networks.
@@
AgentVerse is a multi agent system to study group dynamics (chen2023agentverse). Inspired by human group dynamics, it studies whether a group of expert agents can be more than the sum of its parts. Experiments on text understanding, reasoning, coding, tool utilization, and embodied AI confirm the effectiveness. Problem solving is split into four stages: (1) expert recruitment, (2) collaborative decision making, (3) action execution, and (4) evaluation, where, if the current state is unsatisfactory, a new iteration of the process is started for refinement (see Figure 28). Interestingly, agents manifest emergent behaviors such as volunteering, characterized by agents offering assistance to peers, or conformity, where agents adjust deviated behaviors to align with the common goal under the critics from others. Destructive behaviors were also observed, occasionally leading to undesired and detrimental outcomes.
@@
OASIS is a scalable social media simulator for Twitter/X and Reddit (yang2024oasis). It supports modeling of up to one million LLM-agents. It is built on CAMEL and has role-based agents as its starting point. However, at its large scale, OASIS shows various social group phenomena, including spreading of (mis)information, group polarization, and herd effects. OASIS is built upon an Environment Server, Recommender System, Agent Module, Time Engine, and Scalable Inferencer (see Figure 29).
@@
Research predating LLMs already shows that agent societies may create an automated curriculum of ever increasing difficulty (elman1991incremental; bengio2009curriculum; silver2017mastering; soviany2022curriculum), requiring increasing levels of intelligent behavior from the agents (racaniere2019automated). Similar results have been observed for LLMs (feng2023citing). WebArena is an environment developed to study self-evolving curricula (qi2024webrl), which can also help robot training (ryu2024curricullm) or to mitigate hallucination (zhao2024automatic).
4.3.2. Emergent Social Norms
@@
Social norms play an important role in the predictability of individuals in groups (axelrod1981emergence; axelrod1986evolutionary). Cultural evolution studies how norms evolve at a society level when individuals transmit behavior through imitation, communication, and education (boyd1988culture). LLMs endow agents with the ability to communicate in natural language and have created more opportunities for multi-agent research into societies and the emergence of conventions and norms. Extensive overviews of such new possibilities are provided in (mou2024individual; savarimuthu2024harnessing; xi2023rise). We discuss some of the new approaches in more detail.
@@
EvolutionaryAgent (li2024agent) studies agent alignment in a multi-agent system, with evolutionary methods that go beyond Reinforcement Learning from Human Feedback (see Figure 30). In the context of agent alignment to norms, the approach is controlled: it does not permit the evolution of social norms to be disorderly or random, but it also does not intervene in each step of their evolution. The authors define the initial social norms and a desired direction of evolution. Agents with higher fitness (more norm-conforming) are more likely to reproduce, leading to the diffusion of their strategies, gradually stabilizing and forming new social norms.
@@
Defining a complete, realistic, and complex virtual society is challenging. The purpose of the work is to study how, if such a virtual society existed, a system could further enable evolving intricate evolutionary behaviors of agents, and how this would lead to the emergence of new dynamics. The system provides a sandbox for investigating the safety of AI systems before they impact the real world.
@@
A different approach is based on steels1995self’s naming game, implemented with agents powered by LLaMa 3 and Claude 3.5 (kouwenhoven2024searching; ashery2024dynamics; baronchelli2023shaping). They find that globally accepted conventions or norms can spontaneously arise from local interactions between communicating LLMs. The authors also demonstrate how strong collective biases can emerge during this process, even when individual agents appear to be unbiased, and how minority groups of committed LLMs can drive social change by establishing new social conventions that can overturn established behaviors.
@@
The emergence of norms is studied at another level by horiguchi2024evolution. They explore the potential for LLM agents to spontaneously generate and adhere to normative strategies, building upon the foundational work of Axelrod’s metanorm games. Metanorms are norms enforcing the punishment of those who do not punish agents that are breaking norms (axelrod1986evolutionary). Controlling for personality traits vengefulness and boldness, they find that through dialogue, LLM agents can form complex social norms, metanorms, purely through natural language interaction. A related study evaluates the capability of LLMs to detect norm violations (he2024norm). Based on simulated data from 80 stories in a household context, they investigated whether 10 norms are violated, and found ChatGPT-4 being able for detect norm violations, with Mistral some distance behind.
@@
qiu2024evaluating go a step beyond norms, and study the cultural and social awareness of LLM agents. They introduce CASA, a benchmark designed to assess LLM agents’ sensitivity to cultural and social norms across two web-based tasks: online shopping and social discussion forums. (CASA is based on WebArena (qi2024webrl).) Current LLMs perform significantly better in non-agent than in web-based agent environments, with agents achieving less than 10% awareness coverage and over 40% violation rates. However, using prompting and finetuning on specific datasets, cultural and social awareness can be improved.
@@
Inspired by Society of Mind (minsky1988society), cooperation mechanisms are explored in zhang2023exploring’s agentic LLM simulation. This simulation consists of four unique societies of LLM agents, where each agent is characterized by a specific trait (easy-going or overconfident) and engages in cooperation with a distinct thinking pattern (debate or reflection). They find that LLM agents show human-like social behaviors, such as conformity and consensus reaching, mirroring foundational social psychology theories. Figure 31 shows societies with different types of agents.
@@
The question whether groups of LLM agents can successfully engage in cross-national collaboration and debate is studied by (baltaji2024conformity). Multi agent discussions can support collective AI decisions that reflect diverse perspectives, although agents are susceptible to conformity due to perceived peer pressure. They can also lose track of their personas and opinions. Instructions that encourage debate increase the risk of errors.
4.3.3. Open-World Agents
@@
An important driver of agentic LLM research is the problem of plateauing LLM performance due to limited training data. Open World multi-agent interaction aims to address this problem, generating new interaction data with multi agent simulations. Machine learning can learn no more complexity than what is present in the dataset (or environment). The idea of an open world-model is that it can create infinite datasets or environments, in which agents can continue to learn, to keep improving their intelligence. How should such unlimited challenges be created? The advent of LLMs has given a new impulse to this research question: LLMs are used to solve an LLM-generated problem. This idea is followed, for example, in the multi-agent finetuning approach (subramaniam2025multiagent).
@@
Current agents are primarily created and tested in simplified synthetic environments, leading to a disconnect with real-world scenarios. zhou2023webarena build an environment, Webarena, that is more realistic and reproducible. WebArena is an environment with fully functional websites from four common domains: e-commerce, social forum discussions, collaborative software development, and content management.
@@
Games are eminently suited as open-ended benchmarks for interactive behavior. Real-world tasks require handling intricate interactions, advanced spatial reasoning, long-term planning, and continuous exploration of new strategies. Balrog (paglieri2024balrog) incorporates reinforcement learning environments of varying levels of difficulty, ranging from tasks that are solvable by non-expert humans in seconds to challenging ones that may take years to master (such as the NetHack Learning Environment). They find that while current models achieve partial success in the easier games, they struggle significantly with more challenging tasks such as vision-based decision-making.
@@
Progress in machine learning depends on benchmark availability. As models evolve, there is a need to create benchmarks that can measure progress on new generative capabilities (butt2024benchagents). BenchAgents decomposes the benchmark creation process into planning, generation, data verification, and evaluation, each of which is executed by an LLM agent. These agents interact with each other and utilize human-in-the-loop feedback to explicitly improve and flexibly control data diversity and quality. BenchAgents creates benchmarks to evaluate capabilities related to planning and constraint satisfaction.
@@
AgentBoard provides an evaluation of the breadth of existing benchmarks (ma2024agentboard). Benchmarks should have task diversity. It is necessary to cover various agent tasks such as embodied, web, and tool tasks. Additionally, multi round interaction is important, to mimic realistic scenarios. Existing benchmarks typically adopt single-round tasks. Furthermore, agents should be evaluated in partially-observable environments, to test if they can actively explore their surroundings. Existing agent benchmarks fail to satisfy all of these criteria (ma2024agentboard).
4.4. Discussion
@@ 4.4.1. Interaction Studies
When interacting in multi-agent environments, agentic LLMs show varying levels of performance on games with strategic and coordinated behavior. Enhancing models with reasoning capacities boosts performance, as is evidenced by GPT-o1’s better overall performance and the positive effect of adding probabilistic graphs (xu2024magic). Pre-defined roles and interaction protocols (cooperative or adversarial) help structure the communication between LLM agents while improving task performance. Role-playing frameworks, AI feedback loops, and debate moderation suggests that carefully coordinating multiple LLMs can harness their collective intelligence and yield outcomes that surpass single model performance.
@@
4.4.2. In Depth: CAMEL and Generative Agents
We have surveyed the individual methods for social interaction. In order to dig deeper and highlight some of the issues in this area, we will now discuss two approaches, CAMEL, and Generative Agents, in more detail.
CAMEL
In Section 4.2.2 we saw how LLMs can be prompted to perform different roles, and work together in solving tasks. CAMEL (li2023camel) is a multi-agent system that has been designed to perform this role-playing-for-problem-solving task. Just like the assistants in Section 3.3, the goal is to solve a task, for example in medicine, finance, or computer programming (see also Figure 22). One approach is to write a single monolithic prompt for the LLM in which all the instructions to solve the task are specified—or so the author of the prompt hopes. To measure the success of this approach, a benchmark of test-cases can be created that the LLM has to solve.
The approach of CAMEL is different. CAMEL starts with the idea that specialized agents, with different prompts, may work better. Furthermore, the idea of CAMEL is that the different agents may work better together, and that new problem solving approaches may emerge, that were not present in the single monolithic prompt, or in the initial prompts of the individual agents. CAMEL is a multi-agent system that consists of two agents, an assistant and a user. The user has the domain knowledge, can provide a task specification, as well as feedback on intermediate deliverables that the assistant makes. The assistant has certain skills, for example Python coding, to write a program to solve the problem that the user agent specifies. Both agents are given a so-called inception prompt, an initial idea. They then further send messages (prompts) to each other as they work towards solving the task.
The CAMEL paper describes an experiment where a powerful LLM, GPT4, generates the initial prompt, and a cheaper LLM, GPT-3.5-turbo, executes the further steps. Ten different tasks have been simulated by a combination of 50 agents and 50 users, for a total of 25000 conversations (that were depth-limited to 40 messages).
In such a setup where agents work with agents, the training process may become unstable. CAMEL reports problems where the assistant simply repeats instructions (instead of answering them), problems with fake replies, empty replies, infinite loops of messages, and role reversal. A special (third) critic agent was introduced to ensure a constructive communication process between the agents.
The final experiments report success. The multi-agent system performed better than a single prompt, with performance on the HumanEval benchmark improving form 30% to around 50%. The CAMEL experiments showed that a multi-agent LLM system can be used for problem solving, and the authors showed how stable learning can be achieved when attention is paid to a constructive communication process.
Generative Agents
The architecture of CAMEL recognized two types of agents, assistant and user. The next two papers that we discuss, park2023generative; park2024generative, are about simulating a society with a larger number of agents, from 10-1000. The first paper (park2023generative) introduces the generative agents’ architecture, an architecture that was designed with the aim of generating believable proxies of human behavior. The architecture consists of a memory, a planning module, and a reflection module. In this design agents are given roles and operate in a Sims-like environment. An example is described where they decide to organize a Valentine’s day party (see also Figure 27, and the live web simulation).555See https://reverie.herokuapp.com/arXiv_Demo/ The focus of this work is on achieving believable social behavior, that is unscripted: behavior that emerges from the communications by the agents. We quote the paper: agents wake up, cook breakfast, and head to work; artists paint, while authors write; they form opinions, notice each other, and initiate conversations; they remember and reflect on days past as they plan the next day. The number of agents is larger than in CAMEL, arond 5-15, and the agent architecture is also more involved. Where in CAMEL the behavior was specified fully in-context, in the prompts that are exchanged, in Generative Agents there are additional external algorithms for memory, planning, and reflection. The paper further notes that: the new goals require architectures that manage constantly-growing memories as new interactions, conflicts, and events arise and fade over time while handling cascading social dynamics that unfold between multiple agents, and that success requires an approach that can retrieve relevant events and interactions over a long period, reflect on those memories to generalize and draw higher-level inferences, and apply that reasoning to create plans and reactions that make sense in the moment and in the longer-term of the agent’s behavior. The LLM that is used is GPT-3.5, the same as in CAMEL.
Whether the Generative Agents system generates believable behavior is evaluated with the help of 100 human evaluators. These were enlisted to rank believability of the communication patterns on four categories, by interviewing the agents to probe their ability to (1) remember past experiences, (2) plan future actions based on their experiences, (3) react appropriately to unexpected events, and (4) reflect on their performance to improve their future actions. The evaluation studied the results of four different agent architectures (full architecture, and no observation, no refection, no planning), and found that the full architecture performed best. They also found that generative agents remember with embellishments, and that reflection is required for synthesis of memories. The evaluation did find evidence of emerging communication, relationship building, and coordination. They also found evidence of erratic behavior, hallucination, and misclassification, such as agents that were trying to enter stores after closing time, not understanding the concept of closing time.
Simulating 1000 People
We will now turn to the second paper (park2024generative). Multi-agent simulations can be used to study different aspects of emergent individual and social human behavior. An important methodological challenge in prompt-based simulation studies is to determine how much of the behavior is scripted, and how much emerges. park2024generative focus in the second study on how realistic the behavior of synthetic agents can be. The study is based on their earlier work, with LLM agents that have memory and reflection. A group of 1052 human individuals were recruited who were asked to provide a two hour long interview. The interviews were standardized, administered by an AI interviewer. Next, LLM agents were trained on the audio interview, yielding 1052 different LLM agent profiles. The LLM is prompted to replicate individuals’ attitudes and behaviors and generate synthetic agents.
To validate the accuracy of the personality and behavior of these generative synthetic agents, the agents were tested by interviewing them. As a control group, the human subjects were also interviewed, again, two weeks after their initial interview, to control for natural variation between two interview sessions. The generative agents replicate participants’ responses on the General Social Survey 85% as accurately as participants replicate their own answers two weeks later, and perform comparably in predicting personality traits and outcomes in experimental replications. Subsequently, the synthetic agents have been made available for further experiments.
4.4.3. Emergent Collective Behavior